4x-3y+z=-8

-2x+y-3z=-4

x-y+2z=3


1
Expert's answer
2021-12-14T08:09:27-0500

Question is incomplete.

Let us solve for "x,y , z"

"4x-3y+z=-8\\

-2x+y-3z=-4\\

x-y+2z=3"

Step 1: Swap Row 1 and Row 3 . The result is:

"\begin{array}{ccc}

x -y+2 z =3...(1) \\

-2 x+y +2 z =-4...(2) \\

4 x-3 y+z =-8...(3)

\end{array}"

Step 2: Multiply first equation by 2 and add the result to the second equation. The result is:

"\begin{aligned}

x-y+2 z =3 \\

-y-y+2 z =2 \\

4 x-3 y+z =-8

\end{aligned}"

Step 3: Multiply first equation by -4 and add the result to the third equation. The result is:

"\begin{aligned}

x-y+2 z =3 \\

-y+2 z =2 \\

y -7 z=-20

\end{aligned}"

Step 4: Multiply second equation by 1 and add the result to the third equation. The result is:

"\begin{aligned}

x-y+2 z =3 \\

-y+2 z =2 \\

-5 z=-18

\end{aligned}"

Step 5: solve for z.

"\begin{aligned}

-5 z =-18 \\

z =\frac{18}{5}

\end{aligned}"

Step 6: solve for y using second equation:

"\begin{aligned}

-y+2 z =2 \\

-y+2 \cdot \frac{18}{5} =2 \\

y =\frac{26}{5}

\end{aligned}"

Step 7: solve for x by substituting "y=\frac{26}{5}" and "z=\frac{18}{5}" into the first equation.

"x-\frac{26}{5}+2\times\frac{18}{5}=3\\
x-\frac{26}{5}+\frac{36}{5}=3\\
x+2=3\\
x=1"

Hence, "x=1, y=\frac{26}{5}, z =\frac{18}{5}"